# Quantum Algorithms for the Pathwise Lasso

### v2

@@ 5.1 Slow rates
In the classical linear Gaussian model, the design matrix is deterministic, meaning that $\mathbf{X}\in\mathbb{R}^{n\times d}$ is fixed, while the noise vector $\mathbf{w}\in\mathbb{R}^{n}$ has i.i.d. $\mathcal{N}(0,\sigma^{2})$ entries. We then bound the probability that $\lambda\geq\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}/(1-\epsilon)$. Using that $\mathbf{X}_{i}^{\top}\mathbf{w}$ is a sub-Gaussian random variable with parameter $\sigma\|\mathbf{X}_{i}\|_{2}\leq\sigma\sqrt{Cn}$ (Fact 7), we bound the complement of this probability:
@@
In compressed sensing, it is common for the design matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ to be chosen by the user, and a standard choice is the standard Gaussian matrix with i.i.d. $\mathcal{N}(0,1)$ entries. Assume further that the noise vector $\mathbf{w}\in\mathbb{R}^{n}$ is deterministic with $\|\mathbf{w}\|_{\infty}\leq\sigma$. Thus $\mathbf{X}_{i}^{\top}\mathbf{w}$ is a sub-Gaussian random variable with parameter $\|\mathbf{w}\|_{2}\leq\sqrt{n}\|\mathbf{w}\|_{\infty}\leq\sqrt{n}\sigma$ (Fact 7). By following the exact same steps as the previous example (this time with $C=1$), we conclude that, with probability at least $1-\delta$,
@@ 5.2 Fast Rates
$\|\mathbf{X}(\widetilde{\bm{\beta}}-\bm{\beta}^{\ast})\|_{2}^{2}\leq 81\lambda^{2}\frac{|S|}{n\kappa}+18\lambda\epsilon\|\bm{\beta}_{S}^{\ast}\|_{1},\quad\quad\|\widetilde{\bm{\beta}}-\bm{\beta}^{\ast}\|_{1}\leq 27\lambda\frac{|S|}{n\kappa}+6\epsilon\|\bm{\beta}_{S}^{\ast}\|_{1}.$
@@ 6 Discussion and future work
We point out some future directions of research following our work. In Algorithm 3 and Algorithm 4, the crossing times are computed exactly and the joining times are estimated up to some error. A natural extension would be to allow errors also in computing the crossing times. This would require new techniques in order to retain the path solution continuity. Another direction is to reduce the number of iterations of the LARS algorithm. In the classical setting, Mairal and Yu [78] proposed an approximate LARS algorithm with a maximum number of iterations by employing first-order optimisation methods when two kinks are too close to each other. Finally, it would be interesting to design a fully quantum LARS algorithm by using efficient quantum subroutines for matrix multiplication and matrix inversion, e.g., based on block-encoding techniques [61, 44, 16].
@@ 7 Acknowledgements
JFD thanks Hanzhong Liu and Bin Yu for very useful clarifications regarding Ref. [78] and Lasso in general. JFD thanks András Gilyén for Ref. [4], Yihui Quek for Ref. [85], and Rahul Jain, Josep Lumbreras, and Marco Tomamichel for interesting discussions. JFD also thanks Iosif Pinelis for answering a question on MathOverflow [69]. JFD acknowledges funding from ERC grant No. 810115-DYNASNET. DL acknowledges funding from the Latvian Quantum Initiative under EU Recovery and Resilience Facility under project no. 2.3.1.1.i.0/1/22/I/CFLA/001 in the final part of the project. CSP gratefully acknowledges Ministry of Education (MOE), AcRF Tier 2 grant (Reference No: MOE-T2EP20220-0013) for the funding of this research. This research is supported by the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant and its Quantum Engineering Programme under grant NRF2021-QEP2-02-P05.

### v3

@@ 6.1 Slow rates
$\mathcal{L}_{\lambda}(\widetilde{\bm{\beta}})-\mathcal{L}_{\lambda}(\bm{\beta}^{\ast})\leq\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}\implies\frac{1}{2}{\|}\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}{\|}_{2}^{2}+\lambda(1-\epsilon){\|}\widetilde{\bm{\beta}}{\|}_{1}\leq\frac{1}{2}{\|}\mathbf{y}-\mathbf{X}\bm{\beta}^{*}{\|}_{2}^{2}+\lambda{\|}\bm{\beta}^{*}{\|}_{1}.$
@@
In the classical linear Gaussian model, the design matrix is deterministic, meaning that $\mathbf{X}\in\mathbb{R}^{n\times d}$ is fixed, while the noise vector $\mathbf{w}\in\mathbb{R}^{n}$ has i.i.d. $\mathcal{N}(0,\sigma^{2})$ entries. We then bound the probability that $\lambda\geq\|\mathbf{X}^{\top}\mathbf{w}\|_{\infty}/(1-\epsilon)$. Using that $\mathbf{X}_{i}^{\top}\mathbf{w}$ is a sub-Gaussian random variable with parameter $\sigma\|\mathbf{X}_{i}\|_{2}\leq\sigma\sqrt{Cn}$ (Fact 4), we bound the complement of this probability:
@@
In compressed sensing, it is common for the design matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ to be chosen by the user, and a standard choice is the standard Gaussian matrix with i.i.d. $\mathcal{N}(0,1)$ entries. Assume further that the noise vector $\mathbf{w}\in\mathbb{R}^{n}$ is deterministic with $\|\mathbf{w}\|_{\infty}\leq\sigma$. Thus $\mathbf{X}_{i}^{\top}\mathbf{w}$ is a sub-Gaussian random variable with parameter $\|\mathbf{w}\|_{2}\leq\sqrt{n}\|\mathbf{w}\|_{\infty}\leq\sqrt{n}\sigma$ (Fact 4). By following the exact same steps as the previous example (this time with $C=1$), we conclude that, with probability at least $1-\delta$,
@@ 6.2 Fast Rates
$\|\mathbf{X}(\widetilde{\bm{\beta}}-\bm{\beta}^{\ast})\|_{2}^{2}\leq 81\lambda^{2}\frac{|S|}{n\kappa}+18\lambda\epsilon\|\bm{\beta}_{S}^{\ast}\|_{1}\quad\text{and}\quad\|\widetilde{\bm{\beta}}-\bm{\beta}^{\ast}\|_{1}\leq 27\lambda\frac{|S|}{n\kappa}+6\epsilon\|\bm{\beta}_{S}^{\ast}\|_{1}.$
@@ 7 Discussion and future work
We point out some future directions of research following our work. In Algorithms 3 and 4, the crossing times are computed exactly and the joining times are estimated up to some error. A natural extension would be to allow errors also in computing the crossing times. This would require new techniques in order to retain the path solution continuity. Another direction is to reduce the number of iterations of the LARS algorithm. In the classical setting, Mairal and Yu [80] proposed an approximate LARS algorithm with a maximum number of iterations by employing first-order optimisation methods when two kinks are too close to each other. Moreover, it would be interesting to design a fully quantum LARS algorithm by using efficient quantum subroutines for matrix multiplication and matrix inversion, e.g., based on block-encoding techniques [62, 45, 17]. Also, we believe that our quantum query lower bound is not tight and can be improved, specially the dependence on $\epsilon$ (also not tight in the work of Chen and de Wolf [24]).
@@ 8 Acknowledgements
JFD thanks Hanzhong Liu and Bin Yu for very useful clarifications regarding Ref. [80] and Lasso in general, Yanlin Chen and Ronald de Wolf for helpful clarifications regarding Ref. [24], András Gilyén for Ref. [4], Yihui Quek for Ref. [87], and Rahul Jain, Josep Lumbreras, and Marco Tomamichel for interesting discussions. JFD also thanks Iosif Pinelis for answering a question on MathOverflow [70]. JFD acknowledges funding from ERC grant No. 810115-DYNASNET. DL acknowledges funding from the Latvian Quantum Initiative under EU Recovery and Resilience Facility under project No. 2.3.1.1.i.0/1/22/I/CFLA/001 in the final part of the project. CSP gratefully acknowledges Ministry of Education (MOE), AcRF Tier 2 grant (Reference No: MOE-T2EP20220-0013) for the funding of this research. This research is supported by the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant and its Quantum Engineering Programme under grant NRF2021-QEP2-02-P05.
@@ Appendix A Summary of symbols
$\mathcal{L}^{(\lambda)}_{\mathbf{X},\mathbf{y}}(\bm{\beta})$ Lasso function with input $(\mathbf{X},\mathbf{y})$
¯$\overline{\mathcal{L}}^{(\lambda)}_{\mathbf{X},\mathbf{y}}(\bm{\beta})$ Normalised lasso function with input $(\mathbf{X},\mathbf{y})$
¯$\overline{\mathcal{L}}^{(\lambda)}_{\mathcal{D}}(\bm{\beta})$ Expected Lasso function relative to distribution $\mathcal{D}$